# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

### 29 Jan 2024 v2

@@ 1.4. Szabó’s cocycle category
Before we begin our discussion on actions of $\mathrm{C}^{*}$-tensor categories, we recall the case of a twisted action by a second-countable locally compact group $\Gamma$. We shall later identify this, in the case when $\Gamma$ is countable discrete, with actions of the $*{}^{*}$- tensor category $\mathrm{Hilb}(\Gamma)$.
Definition 1.19 (cf. [40, Definition 1.1]).
Let $\Gamma$ be a locally compact group and $A$ be a $\mathrm{C}^{*}$-algebra. A twisted action of $\Gamma$ on $A$ is a pair $(\alpha,\mathfrak{u})$, where $\alpha:\Gamma\to\mathrm{Aut}(A)$ is a point-norm continuous map, and $\mathfrak{u}:\Gamma\times\Gamma\to\mathcal{U}(\mathcal{M}(A))$ is a strictly continuous map satisfying
(1.5) $\alpha_{1}=\mathrm{id}_{A},\quad\mathrm{Ad}(\mathfrak{u}_{s,t})\circ\alpha_{s}\circ\alpha_{t}=\alpha_{st}$
@@
(1.6) $\mathfrak{u}_{s,1}=\mathfrak{u}_{1,s}=1,\quad\mathfrak{u}_{r,st}\alpha_{r}(\mathfrak{u}_{s,t})=\mathfrak{u}_{rs,t}\mathfrak{u}_{r,s}$
for all $r,s,t\in\Gamma$.
@@
Note that the formulae above differ slightly from the definition of a twisted action in [40]. In fact, the sole difference is that our unitary cocycles $\mathfrak{u}_{s,t}$ for $s,t\in\Gamma$ are the adjoints of the cocycles in [40]. The reason for this change of conventions will be discussed in Section 2.
A triple $(A,\alpha,\mathfrak{u})$ as above is called a twisted $\Gamma$-$\mathrm{C}^{*}$-algebra. If the cocycle is trivial i.e. $\mathfrak{u}_{s,t}=\mathfrak{1}$ for all $s,t\in\Gamma$, then $(A,\alpha)$ is said to be a $\Gamma$-$\mathrm{C}^{*}$-algebra.
Definition 1.21 (cf. [40, Definition 1.10]).
Let $(\alpha,\mathfrak{u}):\Gamma\curvearrowright A$ and $(\beta,\mathfrak{v}):\Gamma\curvearrowright B$ be two twisted actions on $\mathrm{C}^{*}$-algebras $A$ and $B$ respectively.
@@
    A cocycle representation from $(A,\alpha,\mathfrak{u})$ to $(B,\beta,\mathfrak{v})$ is a pair $(\psi,\mathbbm{v})$, where $\psi:A\to\mathcal{M}(B)$ is an extendible $*{}^{*}$-homomorphism and $\mathbbm{v}:\Gamma\to\mathcal{U}(\mathcal{M}(B))$ is a strictly continuous map such that
@@
Due to our change of conventions when defining twisted actions we need to change the definition of a cocycle representation. The formula in (1.8) differs from [40, Definition 1.10] precisely by taking adjoints.
As shown in [40], there exists a category with objects being twisted $\Gamma$-$\mathrm{C}^{*}$-algebras and morphisms being cocycle morphisms with composition defined in [40, Proposition 1.15]. This category is denoted by $\mathrm{C}^{*}_{\Gamma,t}$ ([40, Definition 1.16]).
@@ 2. Actions of $\mathrm{C}^{*}$-tensor categories
For $\mathrm{C}^{*}$-tensor categories $(\mathcal{C},\otimes)$ and $(\mathcal{D},\boxtimes)$, $F:\mathcal{C}\to\mathcal{D}$ is said to be a $\mathrm{C}^{*}$-tensor functor if it is a $\mathrm{C}^{*}$-functor such that $F(1_{\mathcal{C}})=1_{\mathcal{D}}$ there exists a unitary natural isomorphism $J_{X,Y}:F(X)\boxtimes F(Y)\to F(X\otimes Y)$ such that
[…]
@@
In the following definition we denote by $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ the full subcategory of $\mathrm{Corr}_{0}(A)$ consisting of bimodules with countable dense subsets.
@@
[…]
@@
In the literature, the main interest is actions of unitary tensor categories on $*{}^{*}$-algebras (see e.g. [1, 6, 23, 25, 42, 33]). This is because unitary tensor categories axiomatise the standard invariant in subfactor theory and can be thought of as the mathematical objects encoding the symmetry in finite index inclusions of $*{}^{*}$-algebras. If $\mathcal{C}$ is a unitary tensor category, $A$ is separable, and $(F,J):\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ is a $*{}^{*}$-tensor functor, then the bimodule associated to the correspondence $F(X)$ for any $X\in\mathcal{C}$ has a countable dense subset by [25, Corollary 2.24] and hence $(F,J)$ automatically falls into $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$.
@@
Let $\Gamma$ be a a countable discrete group and let $(\alpha,\mathfrak{u}):\Gamma\rightarrow\mathrm{Aut}(A)$ be a twisted action. The pair $(\alpha,\mathfrak{u})$ will induce a $*{}^{*}$-tensor functor $(\alpha,\mathfrak{u}):\mathrm{Hilb}(\Gamma)\rightarrow\mathrm{Corr}_{0}(A)$ by setting
@@
(2.3) $\mathfrak{u}_{\mathbb{C}_{g},\mathbb{C}_{h}}(a\boxtimes b)=\mathfrak{u}_{g,h}^{*}\alpha_{g}(a)b.$
The functor may then be extended by linearity to all of $\mathrm{Hilb}(\Gamma)$ in a similar manner to [14, Proposition 5.6]. In general, actions of $\mathrm{Hilb}(\Gamma)$ on $A$ correspond to twisted actions of $\Gamma$ on $A\otimes\mathbb{K}$.
If $A$ is a $\mathrm{C}^{*}$-algebra, then its sequence algebra $A_{\infty}$ is defined by
@@
[…]
For any $X\in\mathcal{C}$, we can view $F(X)$ as a non-degenerate Hilbert $A$-bimodule. Define
@@
which converges to 00 as $n\to\infty$. Exactly the same calculation shows that if $(\xi_{n})_{n\geq 1}$ and $(\xi_{n}^{\prime})_{n\geq 1}$ induce the same element in $F_{\infty}(X)$ and $(a_{n})_{n\geq 1}\in A_{\infty}$, then $(\xi_{n}\lhd a_{n})_{n\geq 1}=(\xi_{n}^{\prime}\lhd a_{n})_{n\geq 1}$ as elements in $F_{\infty}(X)$. Thus, (2.5) gives a well-defined right action of $A_{\infty}$.
@@
which converges to 00 by Cauchy-Schwarz. Moreover, the sequences $\big{(}\langle\xi_{n},\eta_{n}\rangle_{A}\big{)}_{n\geq 1}$ and $\big{(}\langle\xi_{n}^{\prime},\eta_{n}^{\prime}\rangle_{A}\big{)}_{n\geq 1}$ are bounded, so induce elements in $A_{\infty}$.
Therefore, the function $\langle\cdot,\cdot\rangle_{A_{\infty}}:F_{\infty}(X)\times F_{\infty}(X)\to A_{\infty}$ is well-defined. It is now straightforward to check that this function is right linear, left conjugate linear, and antisymmetric since all this properties are satisfied pointwise for each $n\in\mathbb{N}$. Finally, it is clear that $\langle\xi,\xi\rangle_{A_{\infty}}\geq 0$ and that $\langle\xi,\xi\rangle_{A_{\infty}}=0$ if and only if $\langle\xi_{n},\xi_{n}\rangle_{A}$ converges to 00 i.e. $\xi=0$ in $F_{\infty}(X)$. Since, $F_{\infty}(X)$ is complete with respect to the inner product defined in (2.6), it is a right-Hilbert $A_{\infty}$-bimodule. ∎
@@
As in the proof of Lemma 2.4 we have that (2.5) gives a well-defined left action by $A_{\infty}$. Since the left action of $A$ on $F(X)$ is adjointable, the left action of $(a_{n})_{n\geq 1}\in A_{\infty}$ is an adjointable operator with adjoint given by the action of $(a_{n}^{*})_{n\geq 1}$. Moreover, $F_{\infty}(X)$ is non-degenerate for any $X\in\mathcal{C}$ as $F(X)$ is. ∎
Consider the functor $F_{\infty}:\mathcal{C}\to\mathrm{Corr}_{0}(A_{\infty})$ defined by sending any $X\in\mathcal{C}$ to the correspondence induced by the Hilbert $A_{\infty}$-bimodule $F_{\infty}(X)$, and any $f\in\mathrm{Hom}(X,Y)$ for $X,Y\in\mathcal{C}$ to the intertwiner defined on $F_{\infty}(X)$ by $F_{\infty}(f)((\xi_{n})_{n\geq 1})=(F(f)(\xi_{n}))_{n\geq 1}$. Also, one can define $J_{X,Y}^{\infty}:F_{\infty}(Y)\boxtimes F_{\infty}(X)\to F_{\infty}(X\otimes Y)$ as the unique continuous extension of the map
$J_{X,Y}^{\infty}(\xi\boxtimes\eta)=J_{X,Y}(\xi_{n}\boxtimes\eta_{n}),$
@@
Let $F:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ be an action of a $*{}^{*}$-tensor category $\mathcal{C}$ on a unital $*{}^{*}$-algebra $A$ such that $F(X)$ is algebraically finitely generated as a right $A$-module for all $X\in\mathcal{C}$. Then the map $F_{\infty}:\mathcal{C}\to\mathrm{Corr}_{0}(A_{\infty})$ defined above is a $C^{*}$-tensor functor and so $\mathcal{C}$ acts on $A_{\infty}$ via the triple $(A_{\infty},F_{\infty},J^{\infty})$.
@@
It follows from construction that $F_{\infty}$ is a $*{}^{*}$-functor. Moreover, the naturality of $J^{\infty}$ and that $J^{\infty}$ satisfies commuting diagrams as in (2.1) follows from direct computations. It remains to show that $J_{X,Y}^{\infty}$ is surjective. Let $\iota_{X,Y}:F_{\infty}(Y)\boxtimes F_{\infty}(X)\rightarrow(F(Y)\boxtimes F(X))_{\infty}$ be the canonical inclusion maps and $T_{X,Y}:(F(Y)\boxtimes F(X))_{\infty}\rightarrow F_{\infty}(X\otimes Y)$ be the bounded bimodule map defined by $T(\xi)_{n}=J_{X,Y}(\xi_{n})$ for $\xi=(\xi_{n})_{n\geq 0}\in(F(Y)\boxtimes F(X))_{\infty}$. The composition $T_{X,Y}\circ\iota_{X,Y}$ coincides with $J_{X,Y}^{\infty}$ so it suffices to show the surjectivity of both $T_{X,Y}$ and $\iota_{X,Y}$.
Firstly $T_{X,Y}$ is an adjointable unitary with adjoint defined by the mapping $(\xi_{n})_{n\geq 0}\mapsto(J_{X,Y}^{*}(\xi_{n}))_{n\geq 0}$ for $(\xi_{n})_{n\geq 0}\in F_{\infty}(X\otimes Y)$. It remains to show the surjectivity of $\iota_{X,Y}$. As each $F(X)$ is algebraically finitely generated, then $F(X)$ is also projective ([44, Corollary 15.4.8]). Equivalently for each $X\in\mathcal{C}$ the bimodule $F(X)$ has a right $A$-basis, i.e. there exist a finite set $I_{X}$ and $u_{i}^{X}\in F(X)$ for $i\in I_{X}$ such that $\sum_{i\in I_{X}}u_{i}^{X}\lhd\langle u_{i}^{X},\xi\rangle_{A}=\xi$ for all $\xi\in F(X)$ (see for example [26, Lemma 1.11]). Similarly $(F(Y)\boxtimes F(X))_{\infty}$ is also finitely generated projective as a right $A_{\infty}$-module with right $A_{\infty}$-basis $u_{i}^{Y}\boxtimes u_{j}^{X}\in(F(Y)\boxtimes F(X))_{\infty}$ for $i\in I_{Y}$ and $j\in I_{X}$. As $\iota_{X,Y}(u_{i}^{Y}\boxtimes u_{j}^{X})=u_{i}^{Y}\boxtimes u_{j}^{X}$ and $u_{i}^{Y}\boxtimes u_{j}^{X}$ algebraically generate $(F(Y)\boxtimes F(X))_{\infty}$ as a right $A_{\infty}$-module it follows that $\iota_{X,Y}$ is surjective. ∎
@@
If $F:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ is an action of a unitary tensor category (in the sense of [6] for example) on a unital $*{}^{*}$-algebra $A$, then it follows from [25, 26] that $F(X)$ is algebraically finitely generated as a right $A$-module for each $X\in\mathcal{C}$.

### 08 Apr 2026 v3

@@ 1.4. Szabó’s cocycle category
Before we begin our discussion on actions of $\mathrm{C}^{*}$-tensor categories, we recall the case of a twisted action by a second-countable locally compact group $\Gamma$. We shall later identify this, in the case when $\Gamma$ is countable discrete, with actions of the C- tensor category $\mathrm{Hilb}(\Gamma)$.
Definition 1.17 (cf. [41, Definition 1.1]).
Let $\Gamma$ be a locally compact group and AA be a $\mathrm{C}^{*}$-algebra. A twisted action of $\Gamma$ on AA is a pair $(\alpha,\mathfrak{u})$, where $\alpha:\Gamma\to\mathrm{Aut}(A)$ is a point-norm continuous map, and $\mathfrak{u}:\Gamma\times\Gamma\to\mathcal{U}(\mathcal{M}(A))$ is a strictly continuous map satisfying
(1.5) $\alpha_{1}=\mathrm{id}_{A},\quad\mathrm{Ad}(\mathfrak{u}_{g,h})\circ\alpha_{g}\circ\alpha_{h}=\alpha_{gh}$
@@
(1.6) $\mathfrak{u}_{g,1}=\mathfrak{u}_{1,g}=1,\quad\mathfrak{u}_{k,gh}\alpha_{k}(\mathfrak{u}_{g,h})=\mathfrak{u}_{kg,h}\mathfrak{u}_{k,g}$
for all $k,g,h\in\Gamma$.
@@
Note that the formulae above differ slightly from the definition of a twisted action in [41]. In fact, the sole difference is that our unitary cocycles $\mathfrak{u}_{g,h}$ for $g,h\in\Gamma$ are the adjoints of the cocycles in [41]. The reason for this change of conventions will be discussed in Section 2.
A triple $(A,\alpha,\mathfrak{u})$ as above is called a twisted $\Gamma$-$\mathrm{C}^{*}$-algebra. If the cocycle is trivial i.e. $\mathfrak{u}_{g,h}=\mathfrak{1}$ for all $g,h\in\Gamma$, then $(A,\alpha)$ is said to be a $\Gamma$-$\mathrm{C}^{*}$-algebra.
Definition 1.19 (cf. [41, Definition 1.10]).
Let $(\alpha,\mathfrak{u}):\Gamma\curvearrowright A$ and $(\beta,\mathfrak{v}):\Gamma\curvearrowright B$ be two twisted actions on $\mathrm{C}^{*}$-algebras AA and $B$ respectively.
@@
    A cocycle representation from $(A,\alpha,\mathfrak{u})$ to $(B,\beta,\mathfrak{v})$ is a pair $(\psi,\mathbbm{v})$, where $\psi:A\to\mathcal{M}(B)$ is an extendible -homomorphism and $\mathbbm{v}:\Gamma\to\mathcal{U}(\mathcal{M}(B))$ is a strictly continuous map such that
@@
Due to our change of conventions when defining twisted actions we need to change the definition of a cocycle representation. The formula in (1.8) differs from [41, Definition 1.10] precisely by taking adjoints.
As shown in [41], there exists a category with objects being twisted $\Gamma$-$\mathrm{C}^{*}$-algebras and morphisms being cocycle morphisms with composition defined in [41, Proposition 1.15]. This category is denoted by $\mathrm{C}^{*}_{\Gamma,t}$ ([41, Definition 1.16]).
@@ 2. Actions of $\mathrm{C}^{*}$-tensor categories
For $\mathrm{C}^{*}$-tensor categories $(\mathcal{C},\otimes)$ and $(\mathcal{D},\boxtimes)$, $F:\mathcal{C}\to\mathcal{D}$ is said to be a $\mathrm{C}^{*}$-tensor functor if it is a $\mathrm{C}^{*}$-functor such that $F(1_{\mathcal{C}})=1_{\mathcal{D}}$ and there exists a unitary natural isomorphism $J_{X,Y}:F(X)\boxtimes F(Y)\to F(X\otimes Y)$ such that
(2.1) $(F(X)\boxtimes F(Y))\boxtimes F(Z)$$F(X)\boxtimes(F(Y)\boxtimes F(Z))$$F(X\otimes Y)\boxtimes F(Z)$$F(X)\boxtimes F(Y\otimes Z)$$F((X\otimes Y)\otimes Z)$$F(X\otimes(Y\otimes Z))$$\scriptstyle{\alpha_{F(X),F(Y),F(Z)}}$$\scriptstyle{J_{X,Y}\boxtimes\mathrm{id}_{F(Z)}}$$\scriptstyle{\mathrm{id}_{F(X)}\boxtimes J_{Y,Z}}$$\scriptstyle{J_{X\otimes Y,Z}}$$\scriptstyle{J_{X,Y\otimes Z}}$$\scriptstyle{F(\alpha_{X,Y,Z})}$
@@
In the following definition we denote by $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ the full subcategory of $\mathrm{Corr}_{0}(A)$ consisting of bimodules with countable dense subsets. Moreover, for a $\mathrm{C}^{*}$-tensor category $\mathcal{C}$, we denote by $\mathcal{C}^{\mathrm{rev}}$ the $\mathrm{C}^{*}$-tensor category whose underlying category is $\mathcal{C}$, but the tensor product is reversed i.e. $X\otimes^{\mathrm{rev}}Y=Y\otimes X$.
@@
A $\mathrm{C}^{*}$-tensor category $\mathcal{C}$ is said to act on a $\mathrm{C}^{*}$-algebra AA if there exists a $\mathrm{C}^{*}$-tensor functor $F:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(A)$. If AA is separable we further impose that $F$ is valued in $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$. We will often denote this by $\mathcal{C}\overset{F}{\curvearrowright}A$ or by the triple $(A,F,J)$, where
$J=\{J_{X,Y}:F(Y)\boxtimes F(X)\to F(X\otimes Y)\}_{X,Y\in\mathcal{C}}$
is the natural isomorphism associated with the functor $F$. In this case, we say that the triple $(A,F,J)$ is a $\mathcal{C}$-$\mathrm{C}^{*}$-algebra.
Remark $2.$
Often in the literature an action of $\mathcal{C}$ on AA is given instead by a C-tensor functor $F:\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$. We choose to define it as a functor from $\mathcal{C}^{\mathrm{rev}}$ so that an action of a countable discrete group $\Gamma$ induces an action of the category $\mathrm{Hilb}(\Gamma)$ (see Example 2.4).
@@
In the literature, the main interest is actions of unitary tensor categories on C-algebras (see e.g. [1, 7, 24, 26, 43, 35]). This is because unitary tensor categories axiomatise the standard invariant in subfactor theory and can be thought of as the mathematical objects encoding the symmetry in finite index inclusions of C-algebras. If $\mathcal{C}$ is a unitary tensor category, AA is separable, and $(F,J):\mathcal{C}\rightarrow\mathrm{Corr}_{0}(A)$ is a C-tensor functor, then the bimodule associated to the correspondence $F(X)$ for any $X\in\mathcal{C}$ is of finite index and hence it has a countable dense subset by [26, Corollary 2.24]. Therefore, $(F,J)$ automatically falls into $\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$.
@@
Let $\Gamma$ be a a countable discrete group and let $(\alpha,\mathfrak{u}):\Gamma\rightarrow\mathrm{Aut}(A)$ be a twisted action. The pair $(\alpha,\mathfrak{u})$ will induce a C-tensor functor $(\alpha,\mathfrak{u}):\mathrm{Hilb}(\Gamma)^{\mathrm{rev}}\rightarrow\mathrm{Corr}_{0}(A)$ by setting
@@
(2.3) $\mathfrak{u}_{\mathbb{C}_{g},\mathbb{C}_{h}}(a\boxtimes b)=\mathfrak{u}_{g,h}\alpha_{g}(a)b.$
The functor may then be extended by linearity to all of $\mathrm{Hilb}(\Gamma)^{\mathrm{rev}}$ in a similar manner to [16, Proposition 5.6]. In general, actions of $\mathrm{Hilb}(\Gamma)$ on AA correspond to twisted actions of $\Gamma$ on $A\otimes\mathbb{K}$.
If AA is a $\mathrm{C}^{*}$-algebra, then its sequence algebra $A_{\infty}$ is defined by
@@
We will end this section by showing that if $\mathcal{C}\overset{F}{\curvearrowright}A$ is an action on a separable $\mathrm{C}^{*}$-algebra AA, then we can induce an action of $\mathcal{C}$ on its sequence algebra $A_{\infty}$ whenever the image bimodules are finite rank in the following sense.
Definition 2.5.
We call a right Hilbert module $E$ over a C-algebra AA finite rank if there exist $n\in\mathbb{N}$ and a projection $p\in M_{n}(\mathcal{M}(A))$ such that $E\cong pA^{n}$. In fact, $E$ is finite rank if and only if there exists a collection of elements $\xi_{i}\in\mathcal{L}(A,E)$ for $1\leq i\leq n$ such that
(2.4) $\sum_{i=1}^{n}\xi_{i}\xi_{i}^{*}=\mathrm{id}_{E}.$
We call $\xi_{i}$ a basis for $E$. We say that $(F,J)$ is finite rank if each $F(X)$ for $X\in\mathrm{Irr}(\mathcal{C})$ is finite rank.
Suppose that $F:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}^{\mathrm{sep}}(A)$ is a $\mathrm{C}^{*}$-tensor functor. Our goal is to build a $\mathrm{C}^{*}$-tensor functor $F_{\infty}:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(A_{\infty})$.
For any $X\in\mathcal{C}$, we can view $F(X)$ as a non-degenerate Hilbert AA-bimodule. Define
@@
which converges to 0 as $n\to\infty$. Exactly the same calculation shows that if $(\xi_{n})_{n\geq 1}$ and $(\xi_{n}^{\prime})_{n\geq 1}$ induce the same element in $F_{\infty}(X)$ and $(a_{n})_{n\geq 1}\in A_{\infty}$, then $(\xi_{n}\lhd a_{n})_{n\geq 1}=(\xi_{n}^{\prime}\lhd a_{n})_{n\geq 1}$ as elements in $F_{\infty}(X)$. Thus, (2.6) gives a well-defined right action of $A_{\infty}$.
@@
which converges to 0 by Cauchy-Schwarz. Moreover, the sequences $\big(\langle\xi_{n},\eta_{n}\rangle_{A}\big)_{n\geq 1}$ and $\big(\langle\xi_{n}^{\prime},\eta_{n}^{\prime}\rangle_{A}\big)_{n\geq 1}$ are bounded, so induce the same element in $A_{\infty}$.
Therefore, the function $\langle\cdot,\cdot\rangle_{A_{\infty}}:F_{\infty}(X)\times F_{\infty}(X)\to A_{\infty}$ is well-defined. It is now straightforward to check that this function is right linear, left conjugate linear, and antisymmetric since all these properties are satisfied pointwise for each $n\in\mathbb{N}$. Finally, it is clear that $\langle\xi,\xi\rangle_{A_{\infty}}\geq 0$ and that $\langle\xi,\xi\rangle_{A_{\infty}}=0$ if and only if $\langle\xi_{n},\xi_{n}\rangle_{A}$ converges to 0 i.e. $\xi=0$ in $F_{\infty}(X)$. Since, $F_{\infty}(X)$ is complete with respect to the inner product defined in (2.7) (as a quotient of a complete space by a closed subspace), it is a right-Hilbert $A_{\infty}$-bimodule. ∎
@@
As in the proof of Lemma 2.6 we have that (2.6) gives a well-defined left action by $A_{\infty}$. Since the left action of AA on $F(X)$ is adjointable, the left action of $(a_{n})_{n\geq 1}\in A_{\infty}$ is an adjointable operator with adjoint given by the action of $(a_{n}^{*})_{n\geq 1}$. Moreover, $F_{\infty}(X)$ is non-degenerate for any $X\in\mathcal{C}$ as $F(X)$ is. ∎
Consider the functor $F_{\infty}:\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(A_{\infty})$ defined by sending any $X\in\mathcal{C}$ to the correspondence induced by the Hilbert $A_{\infty}$-bimodule $F_{\infty}(X)$, and any $f\in\mathrm{Hom}(X,Y)$ for $X,Y\in\mathcal{C}$ to the intertwiner defined on $F_{\infty}(X)$ by $F_{\infty}(f)((\xi_{n})_{n\geq 1})=(F(f)(\xi_{n}))_{n\geq 1}$. Also, one can define $J_{X,Y}^{\infty}:F_{\infty}(Y)\boxtimes F_{\infty}(X)\to F_{\infty}(X\otimes Y)$ as the unique continuous extension of the map
$J_{X,Y}^{\infty}(\xi\boxtimes\eta)=(J_{X,Y}(\xi_{n}\boxtimes\eta_{n}))_{n},$
@@
Let $(F,J,A)$ be a finite rank action of a C-tensor category $\mathcal{C}$. Then $\mathcal{C}$ acts on $A_{\infty}$ via the triple $(A_{\infty},F_{\infty},J^{\infty})$.
@@
It follows from construction that $F_{\infty}$ is a C-functor. Moreover, the naturality of $J^{\infty}$ and that $J^{\infty}$ satisfies commuting diagrams as in (2.1) follows from direct computations. It remains to show that $J_{X,Y}^{\infty}$ is surjective. Let $\iota_{X,Y}:F_{\infty}(Y)\boxtimes F_{\infty}(X)\rightarrow(F(Y)\boxtimes F(X))_{\infty}$ be the canonical inclusion maps and $T_{X,Y}:(F(Y)\boxtimes F(X))_{\infty}\rightarrow F_{\infty}(X\otimes Y)$ be the bounded bimodule map defined by $T(\xi)_{n}=J_{X,Y}(\xi_{n})$ for $\xi=(\xi_{n})_{n\geq 0}\in(F(Y)\boxtimes F(X))_{\infty}$. The composition $T_{X,Y}\circ\iota_{X,Y}$ coincides with $J_{X,Y}^{\infty}$ so it suffices to show the surjectivity of both $T_{X,Y}$ and $\iota_{X,Y}$.
Firstly $T_{X,Y}$ is an adjointable unitary with adjoint defined by the mapping $(\xi_{n})_{n\geq 0}\mapsto(J_{X,Y}^{*}(\xi_{n}))_{n\geq 0}$ for $(\xi_{n})_{n\geq 0}\in F_{\infty}(X\otimes Y)$. It remains to show the surjectivity of $\iota_{X,Y}$. Let $X,Y\in\mathcal{C}$ and let $\xi_{i}^{Y}$ with $i\in I_{Y}$ finite such that
$\sum_{i\in I_{Y}}\xi_{i}^{Y}{\xi_{i}^{Y}}^{*}=\mathrm{id}_{F(Y)}.$
Any $\zeta\in(F(Y)\boxtimes F(X))_{\infty}$ is represented by a sequence $(\sum_{l=1}^{k_{n}}y_{l}^{(n)}\boxtimes x_{l}^{(n)})_{n}$. By a standard reindexation one can choose an approximate unit $e_{n}$ for AA such that
${e_{n}\xi_{i}^{Y}}^{*}(y_{l}^{(n)})-{\xi_{i}^{Y}}^{*}(y_{l}^{(n)})\rightarrow 0$
as $n$ tends to infinity for all $i\in I_{Y}$, $1\leq l\leq k_{n}$. Then one has that
$\zeta$ $=\left(\sum_{l=1}^{k_{n}}y_{l}^{(n)}\boxtimes x_{l}^{(n)}\right)_{n}$
$=\left(\sum_{l=1}^{k_{n}}\sum_{i\in I_{Y}}\xi_{i}^{Y}(e_{n}{\xi_{i}^{Y}}^{*}(y_{l}^{(n)}))\boxtimes x_{l}^{(n)}\right)_{n}$
$=\left(\sum_{i\in I_{Y}}\xi_{i}^{Y}(e_{n})\boxtimes\left(\sum_{l=1}^{k_{n}}{\xi_{i}^{Y}}^{*}(y_{l}^{(n)})x_{l}^{(n)}\right)\right)_{n}.$
Hence, it suffices to show that
$\sum_{l=1}^{k_{n}}{\xi_{i}^{Y}}^{*}(y_{l}^{(n)})x_{l}^{(n)}$
is a bounded sequence. This follows precisely as $\zeta$ is a bounded sequence. Indeed,
$\lVert\sum_{l=1}^{k_{n}}{\xi_{i}^{Y}}^{*}(y_{l}^{(n)})x_{l}^{(n)}\rVert^{2}$ $=\|\sum_{l,l^{\prime}}\langle x_{l}^{(n)},({\xi_{i}^{Y}}^{*}(y_{l}^{(n)}))^{*}{\xi_{i}^{Y}}^{*}(y_{l^{\prime}}^{(n)})x_{l^{\prime}}^{(n)}\rangle\|$
$=\|\sum_{l,l^{\prime}}\langle x_{l}^{(n)},({y_{l}^{(n)}}^{*}\xi_{i}^{Y}{\xi_{i}^{Y}}^{*}\boxtimes\mathrm{id}_{F(X)})(y_{l^{\prime}}^{(n)}\boxtimes{x_{l^{\prime}}}^{(n)})\rangle\|$
$\leq\|\sum_{l,l^{\prime}}\langle x_{l}^{(n)},({y_{l}^{(n)}}^{*}\boxtimes\mathrm{id}_{F(X)})(y_{l^{\prime}}^{(n)}\boxtimes{x_{l^{\prime}}}^{(n)})\rangle\|$
$=\|\sum_{l,l^{\prime}}\langle x_{l}^{(n)},\langle{y_{l}^{(n)}},y_{l^{\prime}}^{(n)}\rangle x_{l^{\prime}}^{(n)}\rangle\|$
$=\|\zeta_{n}\|^{2}$
and $\zeta_{n}$ is a bounded sequence. ∎
@@
If $\mathcal{C}\overset{F}{\curvearrowright}A$ is an action of a unitary tensor category (in the sense of [7] for example) on a unital C-algebra AA, then it follows from [26, 27] that $F$ is finite rank.